# Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers

### v2

@@ III Experimental and Simulation Methods
Experimental data was obtained with a home-built confocal microscope as described in [36] and schematically represented in Fig. 2 (c). Single NVs are addressed with a green laser in continuous or pulsed modes using an air objective, while their fluorescence is measured by a single photon detector. MW pulses are produced by an arbitrary waveform generator, which are then amplified and transmitted to the diamond by a thin copper wire. RF fields are generated continuously with controllable power and frequency, combined and then transmitted with the MW pulses as well. The external bias magnetic field $B_{0}$ is applied with a permanent magnet with four degrees of movement, while its intensity and the angle formed with the NV axis $\theta_{0}$ are calculated by the asymmetry between peaks in the continuous wave or pulsed optical detected magnetic resonance (ODMR) spectrum [37]. Lastly, the software package Qudi [38] is used for hardware control and data acquisition.
@@
In order to study each ambiguous resonance, different properties were required for the diamond samples. For measuring 13C ambiguous resonances, Sec. V, a CVD grown sample with 14N was chosen in order to avoid 15N effects. Conversely, in Sec. IV, Hahn echo modulations and XY8-$M$ were measured in shallow implanted 15N diamonds. For longer XY8-12 measurements (Sec. VI), a deeper implanted 15N diamond sample was used with longer coherence times. A more detailed description of the experimental setup and diamond samples can be found in Sec. S1 and Sec. S2.
@@
The simulated time evolutions of the pulsed dynamics were calculated using the master equation solver from Quantum Toolbox in Python (QuTiP) [33, 34] in parallel over 64 physical CPU cores. Experimentally measured fluorescence and simulated transition probability are compared through a linear relation, physically representing experimental uncertainties, as the collection efficiency or background light. The correlation $r$ between them is quantified from -1 to 1 by the Pearson coefficient [1], see Sec. S3 for more details. In Sec. VI, where the coupling parameters of $H_{2}$ are not defined, the transition probability was simulated over several points in parameter space and the simulation with the highest correlation coefficient was assumed.
@@
Figure 3: (a) Experimental and simulated Hahn echo decay for $m_{s}=0\leftrightarrow+1$ transition at $B_{0}=24$ mT and misalignment of $\theta_{0}=2.1^{\circ}$ with modulations due to 15N interaction with NV electronic spin. The coherence decay is fitted by $\exp[-(t/\tau_{c})^{4}]$ [12] with $\tau_{c}=4.018\pm 0.007$ $\mu$s and correlation $r=0.934$. (b) FFT of the simulated Hahn echo decay for both $m_{s}=0\leftrightarrow\pm 1$ transitions. The envelope modulations are defined by the effective Larmor frequencies of 15N $\omega_{m_{s}}^{n}$, numerically simulated from the Hamiltonian. Less intense sidebands from $\omega_{\pm 1}^{n}$ are also present in the spectra. (c) Experimental and simulated XY8-2 of the 15N resonances for different values of $B_{0}$ with $\theta_{0}$=(17.0, 9.3, 5.1, 2.7) and correlations $r$=(0.858, 0.890, 0.971, 0.986), respectively. The resonance peaks have varying amplitude at almost constant pulse separations, evidencing a dominance of the hyperfine interaction with the electronic spin over the Zeeman interaction of the 15N spin with $B_{0}$. The expected resonances from 1H are also shown at $\tau_{H}=1/(2\gamma_{H}B_{0})$. (b) XY8-$M$ at $B_{0}=39$ mT and $\theta_{0}=3.8^{\circ}$ with $r$=(0.975, 0.921, 0.945, 0.906), respectively. At this field, the ambiguous resonance could be mistakenly attributed to 1H nuclear spins. The high correlation between simulations and experimental data for longer pulse sequences further corroborate the numerical model.
IV Magnetic Field Misalignment and 15N Coupling
@@
Perfectly aligning the bias field with the NV axis with accuracy much greater than $\Delta\theta_{0}=1^{\circ}$ requires a precise vector magnet, which is not the case for widely used permanent magnets. In these circumstances, the $S_{x,y}$ Zeeman components and perpendicular nitrogen hyperfine terms can induce undesired electronic spin flips during multipulse sequences. This is better understood in another reference frame where $\vec{B}_{0}$ is aligned along the new $z$ axis, causing the nitrogen hyperfine matrix to assume non-diagonal terms just like in the sensing Hamiltonian of Eq. 3. Consequently, the misaligned field can cause the NV’s electronic spin to sense its own nitrogen nuclear spin. Although this effect happens with both nitrogen isotopes, for 14NV it is nearly negligible at moderate fields due to the quadrupole interaction (Sec. II), thus we restrict the discussion to 15NV in this section.
@@
The influence of 15N coupling in simple pulsed experiments is already known, but restricted to a semi-analytical framework [12, 27] under certain approximations, as discussed in Sec. II. Here we show that the measured Hahn echo decay also agrees with our numerical simulation model with a high correlation of $r=0.934$, Fig. 3 (a). The electronic coherence decay presents envelope modulations with their periods determined by the effective nitrogen Larmor frequencies $\tilde{\omega}_{m_{s}}^{n}$, as shown by the fast Fourier transform (FFT) taken from the simulated data, Fig. 3 (b). As the number of pulses is increased however, an analytical approach for solving Eq. 5 and calculating this nitrogen coupling effect becomes increasingly complex.
@@
To study these effects in longer DD sequences, XY8-2 for the $m_{s}=0\leftrightarrow+1$ transition were measured and simulated over a fixed $\tau$ range varying the magnetic field amplitude and angle, as shown in Fig. 3 (c). The frequency of nitrogen-related resonances does not depend linearly on the applied magnetic field, as one would expect from a simple Zeeman relation. In contrast, the hyperfine interaction, dominating the energy levels of the nuclei, results in peaks with varying intensity and nearly constant position as the field changes. The peaks positions do not correspond to the effective Larmor frequencies of the nitrogen nuclei $\tilde{\omega}_{m_{s}}^{n}$ and a semi-analytical model which describes their resonant frequency values remains an open problem. The lowest magnetic field shows the worst correlation between experiment and simulation of $r=0.858$, where a pulse duration error had to be included in the simulation. This causes a fast modulation of the signal, as discussed in details in . The smaller correlation can be attributed to a larger misalignment angle of $\theta_{0}=17.0^{\circ}$, thus with a larger uncertainty.
@@
These resonances differ in each $m_{s}=0\leftrightarrow\pm 1$ transition due to distinct nuclear state energies and are also present at higher $\tau$ given by the odd fractions of their resonant frequencies ($f_{0}/3$, $f_{0}/5$, $f_{0}/7$…), as previously observed for 13C resonances [6]. At long pulse separations of several $\mu$s, they mix and become a near continuum, which can be visualized in the simulations dataset. 15N resonances do not seem to occur at short pulse spacings of about less than 250 ns. Still, short $\tau$ are occupied by high frequency noises from paramagnetic impurities and even multiples of the resonant frequencies (Sec. VI) [16]. Moreover, setting the resonant frequencies of target nuclear spins with low gyromagnetic ratios $\gamma^{t}$ at $\tau<250$ ns is technically challenging, as it requires large bias fields $B_{0}$ and the capability of amplifying MW fields at a large frequency range of $\Omega_{MW}$ (Eq. 2). Thus, these resonances from 15N are just unavoidable in many cases, underlying the necessity for a dataset where users can compare their experimental data for disambiguation. It is important to note, however, that not all 15NV centers under misaligned fields will present these ambiguous resonances, as additional interactions might dominate over it.
@@
Carbon nuclear isotope 13C with spin-1/2 is present in most diamonds with 1.1% natural abundance, excluding otherwise isotopically purified crystals with increased costs. 13C sensing and manipulation by multipulse sequences is itself an active topic of research in sight of the potential for long-lived quantum memories [6, 40], but in this case, it can act as another source of ambiguous resonances. Compared to 15N in the previous section, the Hamiltonian becomes more complex, as there are six free parameters from the hyperfine interaction $A^{c}$ in Eq. 3 and off-diagonal terms cannot be assumed null because the reference frame is already being chosen such that $\vec{B}_{0}$ lies on the $xz$ plane and the NV axis is along $z$. Therefore, the initial goal is to determine the hyperfine interaction matrix $A^{c}$ between a NV and a coupled 13C nucleus, such that the numerical simulations model can be validated with experimental data and subsequently extended to an arbitrary 13C coupling or even other spins.
@@
Figure 4: (a) Hahn echo decays for both $m_{s}=0\rightarrow\pm 1$ transitions of a 14NV coupled to a 13C nuclear spin at $B_{0}=40$ mT. Both transitions show a slow modulation from the $\tilde{\omega}_{0}^{c}$ nuclear Larmor frequency at $m_{s}=0$ and fast modulations $\tilde{\omega}_{\pm 1}^{c}$ from the $m_{s}=\pm 1$ states. Correlations between experimental and simulated data are $r=$(0.926, 0.761). (b) Modulation frequencies obtained by FFT from the Hahn echo decays as function of $B_{0}$. $\tilde{\omega}_{0}^{c}$ is closely related to the bare gyromagnetic ratio of 13C due to an almost zero hyperfine interaction at $m_{s}=0$, while $\tilde{\omega}_{\pm 1}^{c}$ also decreases or increases linearly with an offset given by Eq. 6. Based on the experimental values of these modulations, the hyperfine tensor $A^{c}$ was determined by comparing them with the calculated eigenvalues from the total Hamiltonian.
@@
Figure 5: (a) XY8-2 resonances from a 14NV coupled to a 13C nuclear spin as a function of $B_{0}$ for the $m_{s}=0\rightarrow+1$ transition. The calculated hyperfine matrix $A^{c}$ was used to simulate the resonances, with correlations $r=$ (0.987, 0.990, 0.989). As the magnetic field increases, the resonance shifts to larger frequencies, influenced the positive gyromagnetic ratio of 13C. (b) Simulated and experimental XY8-$M$ at $B_{0}=40$ mT, with correlations $r=$ (0.989, 0.977, 0.902, 0.951). As the order of the sequence increases, sidebands become more pronounced and line-widths smaller. The high correlation between simulations and experiments attests the robustness of the numerical model also on higher dimensional Hilbert spaces, $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}\otimes\mathcal{H}_{I^{t}}$.
@@
Similar to the coupling with 15N, the Hahn echo decay of an NV coupled to a 13C nuclear spin is also expected to show periodic oscillations given by the effective Larmor frequencies of the nuclei $\tilde{\omega}_{m_{s}}^{c}$ [12]. Thus, Hahn echo was measured in a single NV for both electronic transitions $m_{s}=0\rightarrow\pm 1$ at three different magnetic fields, in a sample with the 14N isotope to avoid nitrogen modulations as in the previous section. The experimentally measured and later simulated modulations at $B_{0}=40$ mT are shown in Fig. 4 (a), where no coherence decay can be observed at this time scale. The Hahn echo for the two other values of $B_{0}$ can be found in Sec. S4. As with 15N coupling, the Hahn echo decay for $m_{s}=0\rightarrow-1$ transition will present a slow modulation from $\tilde{\omega}_{0}^{c}$ and a fast one from $\tilde{\omega}_{-1}^{c}$, similarly to the $m_{s}=0\rightarrow+1$ transition with $\tilde{\omega}_{+1}^{c}$ begin the fast modulation. The values of the nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$ obtained from FFT are plotted against the bias field in Fig. 4 (b). The $\tilde{\omega}_{0}^{c}$ frequency is closely related to the bare gyromagnetic ratio of 13C $\gamma^{c}=10.705$ MHz/T, because of an almost zero hyperfine interaction at $m_{s}=0$. On the other hand, $\tilde{\omega}_{+1}^{c}$ and $\tilde{\omega}_{-1}^{c}$ respectively increases or decreases by the same amount based on the sign of $m_{s}$ and at $B_{0}=0$ mT, both of them are given by [41]
@@
Finally, several different combinations of the hyperfine elements in Eq. 3 were considered and the eigenvalues of the total Hamiltonian $H_{0}+H_{2}$ calculated numerically. The hyperfine interaction matrix $A^{c}$ with the least square difference from the experimental values of $\tilde{\omega}_{m_{s}}^{c}$ is shown in Sec. S4. Using the obtained $A^{c}$, the Hahn echo decay was simulated [Fig. 4 (a)], resulting in the same experimental modulations with high correlations of $r=0.926$ for $m_{s}=+1$ and $r=0.761$ for $m_{s}=-1$.
@@
Based on the obtained hyperfine interaction matrix, XY8-2 for the $m_{s}=0\rightarrow+1$ transition was simulated and measured at the same bias magnetic fields, as shown in Fig. 5 (a). The numerical simulation model and the calculated $A^{c}$ accurately describe the resonances induced by the 13C nuclear spin. As $B_{0}$ increases, the resonance shifts to shorter pulse separations and faster frequencies. This, in contrast to the coupling with 15N [Fig. 3 (b)], evidences that the Zeeman interaction of the 13C nuclear spin with the external magnetic field is comparable in strength with the hyperfine interaction with the NV. Lastly, Fig. 5 (b) presents XY8-$M$ measured and simulated at $B_{0}=40$ mT. Analogous to the 15N resonance [Fig. 3 (d)], the linewidth becomes narrower and sidebands become more pronounced as the order of the sequence increases, which can be understood in terms of an alteration of the frequency filter function induced by the DD sequence [31].
@@
At the magnetic fields shown here, this NV exhibits 13C resonances which do not overlap with other possible nuclear spin species. However, for different couplings and fields, it may be the case that the interaction with 13C leads to ambiguous resonances with other nuclear species. Given the crystalline periodicity of the host diamond, a 13C atom can only occupy a limited number of lattice sites around the NV, resulting in a discrete set of possible hyperfine coupling strengths divided into families [42, 43, 41]. Thus, the dataset provides simulations for hyperfine parameter values calculated from density functional theory [41], such that a few possible 13C resonances can be disambiguated from experimental data at different fields. A more conclusive method to identify the nuclear species giving rise to some resonance is, however, to measure the modulation frequencies on the Hahn echo signal and compare the slow modulation with the expected gyromagnetic ratio, as performed here. Although we focused the discussion on lattice 13C nuclear spins, the same method is applicable to an arbitrary spin coupled to the NV, as long as their interaction can be modeled by Eq. 3.
VI Finite Pulse Duration

### v3

@@ III Experimental and Simulation Methods
Experimental data was obtained with a home-built confocal microscope as described in [35, 36] and schematically represented in Fig. 1 (c). Single NVs are addressed with a green laser in continuous or pulsed modes, while their fluorescence is measured by a single photon detector. MW pulses are produced by an arbitrary waveform generator, which are then amplified and transmitted to the diamond by a thin copper wire. RF fields are generated continuously with controllable power and frequency, combined and then transmitted without amplification with the already amplified MW pulses. The external bias magnetic field $B_{0}$ is applied with a permanent magnet with four degrees of movement, while its intensity and the angle formed with the NV axis $\theta_{0}$ are calculated by the asymmetry between peaks in the continuous wave or pulsed optical detected magnetic resonance (ODMR) spectrum [37]. Lastly, the software package Qudi [38] is used for hardware control and data acquisition.
@@
In order to study each ambiguous resonance, different properties were required for the diamond samples. For measuring 13C ambiguous resonances, Sec. IV.2, a CVD grown sample with 14N was chosen in order to avoid 15N effects. Conversely, in Sec. IV.1, Hahn echo modulations and XY8-$M$ were measured in shallow implanted 15N diamonds. For longer XY8-12 measurements (Sec. IV.3), a deeper implanted 15N diamond sample was used with longer coherence times. A more detailed description of the experimental setup and diamond samples can be found in Appendix A and B.
@@
The time evolution of the pulsed dynamics were calculated using Quantum Color Centers Analysis Toolbox (QuaCCAToo) [39] Python package, based on the master equation solver from Quantum Toolbox in Python (QuTiP) [32, 33]. Some of the most important simulations are open-source and provided in the Example Notebooks section of the QuCAAToo package. Experimentally measured fluorescence and simulated transition probability are compared through a linear relation, physically representing experimental uncertainties, as the collection efficiency or background light. The correlation $r$ between them is quantified from -1 to 1 by the Pearson coefficient [40], see Appendix C for more details. In Sec. IV.3, where the coupling parameters of $H_{2}$ are not defined, the transition probability was simulated over several points in parameter space and the simulation with the highest correlation coefficient was assumed.
@@
Figure 2: (a) Experimental and simulated Hahn echo decay for $m_{s}=0\leftrightarrow+1$ transition at $B_{0}=24$ mT and misalignment of $\theta_{0}=2.1^{\circ}$. The field misalignment leads the Larmor precession of the 15N nuclear spin to be sensed by the electronic spin, resulting in the envelope modulations [41, 24, 26]. The coherence decay is fitted by $\exp[-(t/\tau_{c})^{4}]$ [24] with $\tau_{c}=4.018\pm 0.007$ $\mu$s and correlation $r=0.934$. The y-axis represents the simulated transition probability $p$, while the experimental points are given by the fluorescence counts linearly fitted with the simulation data (Sec. III). This way, some of the experimental points appear above $p=1$ due to experimental noise, even though this would be mathematically impossible. (b) FFT of the simulated Hahn echo decay for both $m_{s}=0\leftrightarrow\pm 1$ transitions. The envelope modulations are defined by the effective Larmor frequencies of 15N $\omega_{m_{s}}^{n}$, numerically simulated from the Hamiltonian. Less intense sidebands from $\omega_{\pm 1}^{n}$ are also present in the spectra. (c) Experimental and simulated XY8-2 of the 15N resonances for different values of $B_{0}$ with $\theta_{0}$=(17.0, 9.3, 5.1, 2.7) and correlations $r$=(0.858, 0.890, 0.971, 0.986), respectively. The resonance peaks have varying amplitude at almost constant pulse separations, evidencing a dominance of the hyperfine interaction with the electronic spin over the Zeeman interaction of the 15N spin with $B_{0}$. The expected resonances from 1H are also shown at $\tau_{H}=1/(2\gamma_{H}B_{0})$. (d) XY8-$M$ at $B_{0}=39$ mT and $\theta_{0}=2.6^{\circ}$ with $r$=(0.975, 0.921, 0.945, 0.906), respectively. At this field, the ambiguous resonance could be mistakenly attributed to 1H nuclear spins. The high correlation between simulations and experimental data for longer pulse sequences further corroborate the numerical model.
IV Results
IV.1 Magnetic Field Misalignment and 15N Coupling
@@
Perfectly aligning the bias field with the NV axis is typically only achievable using permanent magnets with precise rotating stages [26], or superconducting magnets. When considering manually controlled stages, as in our setup, the magnetic field alignment is limited to about $\Delta\theta_{0}=1^{\circ}$. In these circumstances, the terms $S_{x,y}$ in the Zeeman interaction (see Eq. 1.1) and perpendicular nitrogen hyperfine terms can induce undesired electronic spin flips during multipulse sequences. This is better understood in another reference frame where $\vec{B}_{0}$ is aligned along the new $z$ axis, causing the nitrogen hyperfine matrix to assume non-diagonal terms just like in the sensing Hamiltonian of Eq. 3. Consequently, the misaligned field can cause the NV’s electronic spin to sense its own nitrogen nuclear spin. Although this effect happens with both nitrogen isotopes, for 14NV it is nearly negligible at moderate fields due to the quadrupole interaction (Sec. II), thus we restrict the discussion to 15NV in this section.
@@
The influence of 15N coupling in simple pulsed experiments is already well-known [41], resulting in an Electron Spin Echo Envelope Modulation (ESEEM), as first observed by Rowan et al. (1965) [42] in electron paramagnetic resonance experiments. However, these approaches are restricted to a semi-analytical framework [24, 26] under certain approximations, as discussed in Sec. II. Here we show that the measured Hahn echo decay also agrees with our numerical simulation model with a high correlation of $r=0.934$, Fig. 2 (a). The electronic coherence decay presents envelope modulations determined by the effective nitrogen Larmor frequencies $\tilde{\omega}_{m_{s}}^{n}$, as shown by the fast Fourier transform (FFT) taken from the simulated data, Fig. 2 (b). As the number of pulses increases however, an analytical approach for solving Eq. 5 and calculating this nitrogen coupling effect becomes increasingly complex.
@@
To study these effects in longer DD sequences, XY8-2 for the $m_{s}=0\leftrightarrow+1$ transition was measured and simulated over a fixed $\tau$ range varying the magnetic field amplitude and angle, as shown in Fig. 2 (c). The frequency of nitrogen-related resonances does not depend linearly on the applied magnetic field, as one would expect from a simple Zeeman relation. In contrast, the hyperfine interaction, dominating the energy levels of the nuclei, results in peaks with varying intensity and nearly constant position as the field changes. The peaks positions do not correspond to the effective Larmor frequencies of the nitrogen nuclei $\tilde{\omega}_{m_{s}}^{n}$ and a analytical model which describes their resonant frequency values remains an open problem. The lowest magnetic field shows the worst correlation between experiment and simulation of $r=0.858$, where a pulse duration error had to be included in the simulation. This causes a fast modulation of the signal, as discussed in details in Appendix C. This smaller correlation can be attributed to a larger misalignment angle of $\theta_{0}=17.0^{\circ}$, thus with a larger uncertainty.
@@
These resonances differ for each $m_{s}=0\leftrightarrow\pm 1$ transition due to distinct nuclear state energies and are also present at longer $\tau$ values given by the odd multiples of their resonant frequencies, as previously observed for 13C resonances [6]. At long pulse separations of several $\mu$s, they mix and become a near continuum, which can be visualized in the simulations dataset. 15N resonances do not seem to occur at short pulse spacings of about less than 250 ns. Still, short $\tau$ are occupied by high frequency noises from paramagnetic impurities and even fractions of the resonant frequencies (Sec. IV.3) [13]. Moreover, setting the resonant frequencies of target nuclear spins with low gyromagnetic ratios $\gamma^{c}$ at $\tau<250$ ns is technically challenging, as it requires large bias fields $B_{0}$, additionally with broad MW amplifiers for $\Omega_{MW}$ (Eq. 2). Thus, these resonances from 15N are just unavoidable in many cases, underlying the necessity for a dataset where users can compare their experimental data for disambiguation. It is important to note, however, that not all 15NV centers under misaligned fields will present these ambiguous resonances, as additional interactions might dominate over it.
@@
Carbon nuclear isotope 13C with spin-1/2 is present in most diamonds with 1.1% natural abundance. 13C sensing and manipulation by multipulse sequences is itself an active topic of research in sight of the potential for long-lived quantum memories [6, 43] and also for Dynamic Nuclear Polarization applications (DNP) [44, 45], but in this case, it can act as another source of ambiguous resonances. Compared to 15N in the previous section, the Hamiltonian becomes more complex, as there are six free parameters from the hyperfine interaction $A^{c}$ in Eq. 3 and off-diagonal terms cannot be assumed null because the reference frame is already being chosen such that $\vec{B}_{0}$ lies on the $xz$ plane and the NV axis is along $z$. Therefore, the initial goal is to determine the hyperfine interaction matrix $A^{c}$ between an NV and a coupled 13C nucleus, such that the numerical simulations model can be validated with experimental data and subsequently extended to an arbitrary 13C coupling or even other spins.
@@
Figure 3: (a) ESEEM measured with the Hahn echo sequence for both $m_{s}=0\rightarrow\pm 1$ transitions of a 14NV coupled to a 13C nuclear spin at $B_{0}=40$ mT. Both transitions show a slow modulation from the $\tilde{\omega}_{0}^{c}$ nuclear Larmor frequency at $m_{s}=0$ and fast modulations $\tilde{\omega}_{\pm 1}^{c}$ from the $m_{s}=\pm 1$ states. Correlations between experimental and simulated data are $r=$(0.761, 0.926). As in Fig. 2 (a), the experimental fluorescence linearly fitted with the simulated transition probability can reach values above $p>1$ due to experimental noise. (b) Modulation frequencies obtained from the fit of the Hahn echo decay measurements as function of $B_{0}$. $\tilde{\omega}_{0}^{c}$ is closely related to the bare gyromagnetic ratio of 13C (solid line) due to an almost zero hyperfine interaction at $m_{s}=0$, while $\tilde{\omega}_{\pm 1}^{c}$ also decreases or increases by the same amount with an offset given by Eq. 6. Error bars represented by the fit covariance are smaller than the points. Based on the experimental values of these modulations, the hyperfine tensor $A^{c}$ was determined by comparing them with the calculated eigenvalues from the total Hamiltonian.
@@
Figure 4: (a) XY8-2 resonances from a 14NV coupled to a 13C nuclear spin as a function of $B_{0}$ for the $m_{s}=0\rightarrow+1$ transition. The calculated hyperfine matrix $A^{c}$ was used to simulate the resonances, with correlations $r=$ (0.987, 0.990, 0.989). As the magnetic field increases, the resonance shifts to larger frequencies, influenced by the positive gyromagnetic ratio of 13C. (b) Simulated and experimental XY8-$M$ at $B_{0}=40$ mT, with correlations $r=$ (0.989, 0.977, 0.902, 0.951). As the order of the sequence increases, sidebands become more pronounced and line-widths smaller due to changes in the filter function. For this 13C coupling and these fields, these ambiguous resonances do not correspond to other nuclear spin. The high correlation between simulations and experiments attests the robustness of the numerical model also on higher dimensional Hilbert spaces, $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}\otimes\mathcal{H}_{I^{c}}$.
@@
Similar to the coupling with 15N, the Hahn echo decay of an NV coupled to a 13C nuclear spin is also expected to show periodic oscillations given by the effective Larmor frequencies of the nuclei $\tilde{\omega}_{m_{s}}^{c}$ [24]. Thus, Hahn echo was measured in a single 14NV for both electron spin transitions $m_{s}=0\rightarrow\pm 1$ at three different magnetic fields, to avoid nitrogen modulations as in the previous section. The experimentally measured and later simulated modulations at $B_{0}=40$ mT are shown in Fig. 3 (a), where no coherence decay can be observed at this time scale. The ESEEM for the two other values of $B_{0}$ can be found in Appendix D. As with 15N coupling, the Hahn echo decay for $m_{s}=0\rightarrow-1$ transition will present a slow modulation from $\tilde{\omega}_{0}^{c}$ and a fast one from $\tilde{\omega}_{-1}^{c}$, similarly to the $m_{s}=0\rightarrow+1$ transition with $\tilde{\omega}_{+1}^{c}$ begin the fast modulation. The values of the nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$ obtained from the fit of the curve [8] are plotted against the bias field in Fig. 3 (b). The $\tilde{\omega}_{0}^{c}$ frequency is closely related to the bare gyromagnetic ratio of 13C $\gamma^{c}=10.705$ MHz/T, because of an almost zero hyperfine interaction at $m_{s}=0$. On the other hand, $\tilde{\omega}_{+1}^{c}$ and $\tilde{\omega}_{-1}^{c}$ respectively increases or decreases by the same amount based on the sign of $m_{s}$. At $B_{0}=0$ mT, both of them are given by [46]
@@
Finally, several different combinations of the hyperfine elements in Eq. 3 were considered and the eigenvalues of the total Hamiltonian $H_{0}+H_{2}$ calculated numerically. The hyperfine interaction matrix $A^{c}$ with the least square difference from the experimental values of $\tilde{\omega}_{m_{s}}^{c}$ is shown in Appendix D. Using the obtained $A^{c}$, the Hahn echo decay was simulated [Fig. 3 (a)], resulting in the same experimental modulations with high correlations of $r=0.761$ for $m_{s}=+1$ and $r=0.926$ for $m_{s}=-1$.
@@
Based on the obtained hyperfine interaction matrix, XY8-2 measurement for the $m_{s}=0\rightarrow+1$ transition was simulated and measured at the same bias magnetic fields, as shown in Fig. 4 (a). The numerical simulation model and the calculated $A^{c}$ accurately describe the resonances induced by the 13C nuclear spin. As $B_{0}$ increases, the resonance shifts to shorter pulse separations and faster frequencies. This, in contrast to the coupling with 15N [Fig. 2 (c)], evidences that the Zeeman interaction of the 13C nuclear spin with the external magnetic field is comparable in strength with the hyperfine interaction with the NV. Lastly, Fig. 4 (b) presents XY8-$M$ measured and simulated at $B_{0}=40$ mT. Analogous to the 15N resonance [Fig. 2 (d)], the linewidth becomes narrower and sidebands become more pronounced as the order of the sequence increases, which can be understood in terms of an alteration of the frequency filter function induced by the DD sequence [30]. As the order increases, the correlation between experimental and simulation data decreases, possibly due to some pulse-length error [Appendix E] or inaccuracy in the fitted ESEEM frequencies.
@@
At the magnetic fields shown here, this particular NV exhibits 13C resonances which do not overlap with other possible nuclear spin species. However, for different couplings and fields, it may be the case that the interaction with 13C leads to the appearance of ambiguous resonances mimicking other nuclear species. Given the crystalline periodicity of the host diamond, a 13C atom can only occupy a limited number of lattice sites around the NV, resulting in a discrete set of possible hyperfine coupling strengths divided into families [47, 48, 46]. Thus, the dataset provides simulations for hyperfine parameter values calculated from density functional theory [46], such that a few possible 13C resonances can be disambiguated from experimental data at different fields. A more conclusive method to identify the nuclear species giving rise to some resonance is, however, to measure ESEEM frequencies and compare the slow modulation with the expected gyromagnetic ratio, as performed here. Although we focused the discussion on lattice 13C nuclear spins, the same method is applicable to an arbitrary spin coupled to the NV, as long as their interaction can be modeled by Eq. 3.
IV.3 Finite Pulse Duration